OpenStudy (anonymous):

What is the slope and y-intercept of the graph of 3x + 2y = 12?

OpenStudy (anonymous):

Y interceot is 6 slope is -3/2

OpenStudy (precal):

3x+2y=12 y=(12-3x)/2 y=6-(3/2)x m=-3/2 b=6